# Reconfigurable routing in data center networks

### v1

@@ 4 Results
We now present and discuss our results but defer proofs to Section 5. In general terms, we obtain NP-completeness for RRP($\Delta_{S}=2$) and RRP($\Delta_{S}=3$) on any fixed class of static networks of practical interest (defined more fully below) and then restrict our focus (and associated parameters) to the case where the static network is a hypercube, and establish the NP-completeness of RRP($\Delta_{S}=1,\sigma=3$) in this setting.
As is standard in NP-hardness proofs, we reduce from known NP-complete problems to instances of RRP; the challenge is that, due to the expansive scope of our theorems, we lose several “degrees of freedom” which are used for encoding hard instances in e.g. FFS19 ; FFSV20 ; FGS18 . Specifically, we may not make use of varying static or dynamic link weights to prohibit certain connections, nor encode any features of the input instance in the topology of the static network $G$. We are constrained to choose a size for the network $G$, and then to encode the input instance in the demand matrix $D$.
@@
Theorem 8.
For any polynomial family of networks =$\mathcal{H}=\{H_{i}:i\geq 0\}$, the problem RRP$(\Delta_{S}=2,\sigma\geq 0)$ restricted to instances $(H(S),D,\mu,\kappa)$ satisfying:
  • 1.
@@
  • 5.
    $\mu\in\Theta(\frac{1}{poly(n)})$ is fixed for all dynamic links
  • 6.
    $\kappa\in poly(n)$
@@
This result significantly strengthens Theorem 4 as proven in FPS19 : there, RRP($\Delta_{S}\geq 2,\sigma=0$) is shown to be NP-complete when the static network is an independent set, and the proof does not enable us to restrict the workload matrix $D$ meaningfully. The main weakness of Theorem 8 is its reliance on $\mu$ being a polynomial factor smaller than any static link weight. This is actually related to the fact that a connected 2-regular network, as is $G(N)$ when $G$ is an independent set and $\Delta_{S}=2$, has diameter linear in the number of nodes $n$. A network of maximum degree 3, on the other hand, may have diameter logarithmic in $n$ (e.g., a complete binary tree has this property) and we indeed show NP-completeness of RRP($\Delta_{S}=3$) when $\mu=\Theta(\frac{1}{\log n})$.
Theorem 9.
For any polynomial graph family $\mathcal{H}$, the problem RRP$(\Delta_{S}=3,\sigma\geq 0)$ restricted to instances $(H(S),D,\mu,\kappa)$ satisfying:
  • 1.
    $H\in\mathcal{H}$ has size $n$
  • 2.
    $|S|=\{s\}$
  • 3.
    the weight of every static link in $H(S)$ is $1$
  • 4.
    the workload matrix $D$ is sparse and all values in it are polynomial in $n$
  • 5.
    $\mu\in\Theta(\frac{1}{\log(n)})$
  • 6.
    $\kappa\in poly(n)$
is NP-complete.
These results led us to consider the problem RRP($\Delta_{S}=1,\sigma\geq 0$). By extending Theorem 6 of FPS19 , we show that this restriction leads to a tractable problem when either $\sigma=0$ or the static network is a complete graph.
Theorem 10.
RRP($\Delta_{S}=1,\sigma=0$) is in P when there is one switch.
Corollary 1.
RRP($\Delta_{S}=1$) with a single switch to which each node is connected exactly once is in P when the static network is a complete graph with all edges of equal weight.
Corollary 1 rules out the possibility that RRP($\Delta_{S}=1,\sigma>0$) might be NP-complete for any polynomial graph family $\mathcal{H}$ (since such a claim would extend to the family of complete graphs) unless P equals NP. This leaves open the practically relevant case where $\Delta_{S}=1$ and $\sigma>0$ for specific topologies. We consequently consider the scenario where the static network is a hypercube and the segregation parameter $\sigma=3$.
Theorem 11.
For the family of hypercubes $\mathcal{Q}$, the problem RRP($\Delta_{S}=1$, $\sigma=3$) restricted to instances $(H(S),D,\mu,\kappa)$ satisfying:
  • 1.
    $H=Q_{n}$
  • 2.
    $|S|=\{s\}$
  • 3.
    the weight of every static link in $H(S)$ is $1$
  • 4.
    the workload matrix $D$ is sparse and all values in it are polynomial in $n$
  • 5.
    $\mu\in(0,1)$
  • 6.
    $\kappa\in poly(n)$
is NP-complete.
We emphasize the relevance of the hypercube as a prototypical model of interconnection networks (see, e.g., GLL09 ) and the fact that we obtain hardness here for any choice of fixed dynamic link weight $\mu$.
5 Proofs
We now present the proofs of our results given in the previous section.
Theorem 8.
For any polynomial family of networks =$\mathcal{H}=\{H_{i}:i\geq 0\}$, the problem RRP$(\Delta_{S}=2,\sigma\geq 0)$ restricted to instances $(H(S),D,\mu,\kappa)$ satisfying:
  • 1.
    $H\in\mathcal{H}$ has size $n$
  • 2.
    $S=\{s\}$
  • 3.
    the weight of every static link in $H(S)$ is $1$
  • 4.
    the workload matrix $D$ is sparse and all values in it are polynomial in $n$
  • 5.
    $\mu\in\Theta(\frac{1}{poly(n)})$ is fixed for all dynamic links
  • 6.
    $\kappa\in poly(n)$
is NP-complete.
Proof.
Note that RRP as in the statement of the theorem is in NP on account of the numerical restrictions imposed. We now build a polynomial-time reduction from the problem 3-Min-Bisection to our restricted version of RRP where the problem 3-Min-Bisection is defined as follows:
  • 1.
@@ 5 Proofs
Note that any $3$-regular graph necessarily has an even number of nodes. The problem 3-Min-Bisection was proven to be NP-complete in BK02 where it was also shown that approximating 3-Min-Bisection to within a constant factor approximation ratio entails the existence of a constant factor approximation algorithm for the more general and widely-studied problem Min-Bisection, which is defined as above but where $G$ can be arbitrary and where $A$ and $B$ have sizes differing by at most one. Given an arbitrary instance $(G=(V,E),k)$ of size $n$ of 3-Min-Bisection, we now build an instance $(H(S),D,\mu,\kappa)$ of RRP. Moreover, we may assume that $k\leq\frac{n}{3}+46$ as it was proven in CE88 that every 3-regular graph has bisection width at most $\frac{n}{3}+46$.
@@
As stated earlier, our static network $H$ is the network $H_{i}\in\mathcal{H}$ where $|H_{i}|=\bar{n}$. All static links have weight $1$. We refer to the node set of $H$ as $V^{\prime}$ also and we refer to the subset of nodes within $V^{\prime}$ corresponding to $V$ as $V$ also. We have one switch $s$ with $2|V^{\prime}|$ ports so that every node of $H$ is adjacent, via switch links, to exactly two ports of the switch. Hence, our switch set $S=\{s\}$ and our hybrid network is $H(S)$. It is important to note that for any configuration $N$, any node of $H(N)$ can be adjacent to at most $2$ other nodes via dynamic links (as $\Delta_{S}=2$).
@@
Conversely, suppose that $(H(S),D,\mu,\kappa)$ is a yes-instance of RRP. Let $N$ be a configuration (consisting of a switch matching of $s$) and let $F$ be a collection of flow-paths that witness that the total workload cost for the workload matrix $D$ is at most $\kappa$. Denote by $N$, also, the sub-network of $H(N)$ consisting solely of dynamic links. W.l.o.g. we may assume that every node of $H(N)$ is incident with exactly two dynamic links of $N$ (by adding additional dynamic links that we do not actually use in any flow-path, if necessary).
Claim 1.
@@
Suppose that the paths $p_{A}$ and $p_{B}$ have different lengths. From above, the total workload cost of the flow-paths serving the $\beta$-demands is at least $3n^{4}+n(2(\frac{n}{2}+1)^{2}-2n(\frac{n}{2}+1)+n^{2}+n)=3n^{4}+\frac{n^{3}}{2}+n^{2}+2n=\kappa_{\beta}+2n$. Consequently, we must have that $2n\leq\kappa_{1}=\frac{k}{2}+\frac{1}{8}-\frac{1}{4n}+\frac{k}{3n^{2}}$. As $k\leq\frac{n}{3}+46$ (from CE88 ), we have that $2n\leq\frac{n}{6}+23+\frac{1}{8}-\frac{1}{4n}+\frac{1}{9n}+\frac{46}{3n^{2}}$ which yields a contradiction. Hence, the paths $p_{A}$ and $p_{B}$ each have length $\frac{n}{2}$, with the total workload cost of the flow-paths corresponding to the unit-edges being at most $\kappa_{1}$.
@@
Let $A$ (resp. $B$) be the nodes of $V$ that appear on $p_{A}$ (resp. $p_{B}$). So, $(A,B)$ is a bisection of $G$. Suppose that this bisection has width $k+\epsilon$, for some $\epsilon\geq 1$. So there are $k+\epsilon$ unit-edges whose corresponding work-flows contribute collectively at least $(k+\epsilon)\mu(L+2)\geq(k+1)(\frac{1}{2}+\frac{1}{3n^{2}})=\frac{k}{2}+\frac{1}{2}+\frac{1}{3n^{2}}+\frac{k}{3n^{2}}>\frac{k}{2}+\frac{1}{8}-\frac{1}{4n}+\frac{k}{3n^{2}}=\kappa_{1}$ to the total workload cost, which yields a contradiction. Consequently, $G$ has bisection width at most $k$ and we have that if $(H(S),D,\mu,\kappa)$ is a yes-instance of RRP then $(G,k)$ is a yes-instance of 3-Min-Bisection. Our result follows as $(H(S),D,\mu,\kappa)$ can be constructed from $(G,k)$ in time polynomial in $n$.
Theorem 9.
For any polynomial graph family $\mathcal{H}$, the problem RRP$(\Delta_{S}=3,\sigma\geq 0)$ restricted to instances $(H(S),D,\mu,\kappa)$ satisfying:
  • 1.
@@
  • 5.
    $\mu\in\Theta(\frac{1}{\log(n)})$
  • 6.
    $\kappa\in poly(n)$
@@
Proof.
Note that RRP as in the statement of the theorem is in NP on account of the numerical restrictions imposed. The problem RXC3 is defined as follows:
  • 1.
@@
The problem RXC3 is NP-complete as was proven in Theorem A.1 of Gon85 . Let $(\mathcal{X},\mathcal{C})$ be an instance of RXC3 of size $n$. We will build an instance $(H(S),D,\mu,\kappa)$ of RRP corresponding to $(\mathcal{X},\mathcal{C})$. First, we define a weighted graph $D^{\prime}$ so as to describe the workload $D$. Having built $D^{\prime}$, we will orient every undirected edge so that: there is a directed edge $(u,v)$ with weight $w$ if, and only if, there is a node-to-node workload of $w$ from $u$ to $v$ (we still refer to the resulting digraph as $D^{\prime}$). Consequently, we will need one flow-path in our resulting hybrid network corresponding to each directed edge of $D^{\prime}=(V^{\prime},E^{\prime})$. Our construction proceeds in stages.
  • 1.
@@
Let $\bar{n}$ be the size of the network $H_{i}$ where $i$ is the smallest integer such that $N\leq|H_{i}|$ and we fix our static network as $H=H_{i}$. We name a subset of nodes of $H$ as $V^{\prime}$ with the remaining nodes named $U$. We have one switch $s$ that has $3n$ switch ports and every node of $H$ has $3$ external ports; so, every node of $H$ has exactly $3$ switch links to $s$. We set $S=\{s\}$. Finally, we define $\mu=\frac{1}{2\lceil\log_{2}(n)\rceil}$ and $\kappa=\mu\alpha|E^{\prime\prime}|+3\mu n(\lceil\log_{2}(n)\rceil+3)$ so as to complete the construction of our instance $(H(S),D,\mu,\kappa)$ of RRP.
Suppose that $(\mathcal{X},\mathcal{C})$ is a yes-instance of RXC3. So, let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be an exact cover of $\mathcal{X}$. For every $(u,v)\in E^{\prime\prime}$, we choose $(u,v)$ to be a dynamic link and ensure that every node of $\{v_{j}:1\leq j\leq 3n,c_{j}\in\mathcal{C}^{\prime}\}$ is joined via a dynamic link to a leaf node of $\{l_{i}:1\leq i\leq n\}$. If $v_{j}$ is such that $c_{j}\in\mathcal{C}^{\prime}$ then denote the unique leaf node to which there is a dynamic link from $v_{j}$ by $\bar{l}_{j}$. Denote the resulting switch matching of $s$ as constituting the configuration $N$. Referring back to our comment on node degrees above, the dynamic links so added must form a matching in $H(N)$. For any workload of weight $\alpha$ in $D$, we choose the corresponding flow-path to consist of the dynamic link joining the two nodes in question. Consequently, for all of the workloads corresponding to directed edges of $E^{\prime\prime}$, the total workload cost of the corresponding flow-paths is $\mu\alpha|E^{\prime\prime}|$. All of the remaining workloads involve the root and the nodes of $\{x_{i}:1\leq i\leq n\}$. Any one of these nodes $x_{i}$ is such that the element $x_{i}$ in a unique clause $c_{j}\in\mathcal{C}^{\prime}$. We choose the flow-path consisting entirely of dynamic links that routes from the root down the tree $T$ to the leaf-node $\bar{l}_{j}$ and on to $v_{j}$ and, through either $u_{j}$ or $w_{j}$, on to $x_{i}$. This flow-path has (graph-theoretic) length $d+3$ and consequently the total workload cost due to flow-paths corresponding to edges of $E^{root}$ is exactly $3\mu n(d+3)=3\mu n(\lceil\log_{2}(n)\rceil+3)$. Hence, the total workload cost due to all flow-paths is $\mu\alpha|E^{\prime\prime}|+3\mu n(\lceil\log_{2}(n)\rceil+3)=\kappa$ and $(H(S),D,\mu,\kappa)$ is a yes-instance of RRP.
Conversely, suppose that $(H(S),D,\mu,\kappa)$ is a yes-instance of RRP and that $N$ is a configuration and $F$ a set of flow-paths witnessing that the total workload cost is at most $\kappa$.
Claim 1.
@@
Consider the flow-path corresponding to the directed edge $(r,x_{i})\in E^{root}$. Suppose that the dynamic link $(l_{k},v_{j})$ appears on this flow-path (exactly one such dynamic link does). Note that there are at most $n$ such dynamic links used in the flow-paths of $F$ corresponding to the directed edges of $E^{root}$. Build the set of clauses $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ by including every such $c_{j}$ in $\mathcal{C}^{\prime}$. Consequently, we have a subset $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ of at most $n$ clauses. Moreover, as we can reach any $x_{i}$ from $r$ by a flow-path of $d+3$ dynamic links, every $x_{i}\in\mathcal{X}$ appears in a clause of $\mathcal{C}^{\prime}$; that is, $|\mathcal{C}^{\prime}|=n$, $\mathcal{C}^{\prime}$ is an exact cover for $\mathcal{X}$ and $(\mathcal{X},\mathcal{C})$ is a yes-instance of RXC3. As our instance $(H(S),D,\mu,\kappa)$ can be constructed from $(\mathcal{X},\mathcal{C})$ in time polynomial in $n$, our result follows.
Recall Theorem 6. We extend this to obtain two tractability results which usefully contrast with our NP-completeness results.
Theorem 6 ((FPS19, , Theorem 3.1)).
@@
Theorem 10.
RRP($\Delta_{S}=1,\sigma=0$) is in P when there is one switch.
Proof.
Note that this proof holds even if we allow variable weights for static and optic links. Because each node $v$ is connected to the switch exactly once, it is not possible that any vertex is incident to two optic links, and hence impossible for there to be any path consisting of two or more optic links. Consequently setting $\delta=1$ introduces no new constraints; in this scenario a flow is permissible with $\sigma=0$ if and only if it permissible with $\sigma=0$ and $\delta=1$. Hence the input instance of RRP($\Delta_{S}=1,\sigma=0$) is a yes-instance if and only if it is a yes-instance of 1-switched RRP($\Delta_{S}=1,\sigma=0,\delta=1$) and we have tractability from Theorem 6.
Corollary 1.
RRP($\Delta_{S}=1$) with a single switch to which each node is connected exactly once is solvable in polynomial time when $G$ is a complete graph with all edges of equal weight.
Proof.
@@
  • 1.
    a single optic link from $u$ to $v$ with unit weight.
  • 2.
    a single static link from $u$ to $v$ with weight $\mu$.
It follows that setting $\sigma=0$ introduces no new constraints, and then by Theorem 10 we have tractability.
Theorem 11.
For the family of hypercubes $\mathcal{Q}$, the problem RRP($\Delta_{S}=1$, $\sigma=3$) restricted to instances $(H(S),D,\mu,\kappa)$ satisfying:
  • 1.
    $H=Q_{n}$
  • 2.
    $|S|=\{s\}$
  • 3.
    the weight of every static link in $H(S)$ is $1$
  • 4.

### v2

@@ 3 Results
Table 1 shows a summary of hardness results from previous work as well as our three main intractability results. In general terms, we obtain NP-completeness for 2-switched RRP and 3-switched RRP on any fixed class of static networks of practical interest (defined more fully below) and for any value of $\sigma$. We then restrict our focus (and associated parameters) to the case where the static network is a hypercube when we establish the NP-completeness of 1-switched RRP($\sigma=3$) in this setting; we conjecture that a similar construction can be used to establish hardness when $\sigma>3$. We also, in Theorem 3.3, show that 1-switched RRP($\sigma=0$) is solvable in polynomial time. The cases when $\sigma\in\{1,2\}$ remain interesting open problems. Subsection 3.1 is devoted to the case of 1-Switched RRP($\delta=1$) (i.e. any flow-path must contain at most a single dynamic link) which entails at most 2 alternations, and we show this case is NP-complete for hypercubes, grids, and toroidal grids. The proof in that section is intended to facilitate the task of proving hardness for other interesting families of networks.
Result $|S|$ $\mathrm{\Delta}_{R}$ $\sigma$/$\delta$ $D$ link weights notes
[11], Theorem 1 $\mathrm{\Theta}(n)$    (or 1  [^2] By using variable $\mu$ with prohibitively large weights, it is possible to simulate many switches with just one.) $\mathrm{\Theta}(n)$ any $\sigma\geq 2$ sparse, all values 00 or $1$ variable; $w\in[1,100n^{2}]$ $\mu\in[1,100n^{2}]$ Showed inapprox. within $\mathrm{\Omega}(\log n)$
[12], Lemma 1 $\mathrm{\Theta}(n)$ $1$ fixed; $w=\mu=1$ All switches have 3 ports.
[12], Theorem 2 $1$ $G$ has $\mathrm{\Theta}(n)$ components
[13], Theorems 4.1, 4.2 $2$ any $\sigma\geq 0$ dense, values in $\mathrm{poly}(n)$ $G$ is empty; there are no static links
Theorem 3.1 sparse, values in $\mathrm{poly}(n)$ fixed;        $w=1$    $\mu\in\mathrm{\Theta}(\frac{1}{\mathrm{poly}(n)})$ $G\in\mathcal{H}$, where $\mathcal{H}$ is any polynomial family of networks (incl. hypercubes, grids, cycles).
Theorem 3.2 3 fixed;        $w=1$    $\mu\in\mathrm{\Theta}(\frac{1}{\mathrm{log}(n)})$
Theorem 3.4 $1$ $\sigma=3$ fixed; $w=1$ any $\mu\in(0,1)$ $G$ is a hypercube
Theorem 3.8 $\delta=1$ $G$ is a hypercube, grid, or toroidal grid
Table 1: Settings for some pre-existing hardness results for RRP. $|S|$ is the number of switches; $\mathrm{\Delta}_{R}$ is the maximum number of external ports per node; $\sigma$ is the segregation parameter and $\delta$ is the dynamic link limit; $D$ is the workload matrix; $n$ denotes the number of nodes in the instance.
As is standard in NP-hardness proofs, we reduce from known NP-complete problems to instances of RRP; the challenge is that, due to the expansive scope of our theorems, we lose several “degrees of freedom” which are used for encoding hard instances in, e.g., [10, 11, 12]. Specifically, we may not make use of varying static or dynamic link weights to prohibit certain connections, nor encode any features of the input instance in the topology of the hybrid network $G(S)$. For example, in Lemma 1 [12], many small switches with two feasible configurations each are used to encode a truth assignment, and in Theorem 1 of [11] “bad” links are given weights of order $\mathrm{\Theta}(n^{2})$. Neither of these mechanisms can be leveraged to obtain hardness in our setting; in this sense, our hardness results are strictly stronger and also harder to obtain than those from [10, 11, 12]. We are constrained to choose a size for the network $G$, and then to encode the input instance in the demand matrix $D$.
@@
Theorem 3.1
For any polynomial family of networks =$\mathcal{H}=\{H_{i}:i\geq 0\}$, the problem 2-switched RRP restricted to instances $(H,\mu,D,\kappa)$ satisfying:
@@
  • $\mu\in\mathrm{\Theta}(\frac{1}{poly(n)})$ is fixed for all dynamic links
@@
Proof
Note that 2-switched RRP as in the statement of the theorem is in NP as it can be straightforwardly reduced to an equivalent instance of the more general RRP, which is known to be in NP. We now build a polynomial-time reduction from the problem 3-Min-Bisection to our restricted version of RRP where the problem 3-Min-Bisection is defined as follows:
@@
Note that any $3$-regular graph necessarily has an even number of nodes. The problem 3-Min-Bisection was proven to be NP-complete in [5] where it was also shown that approximating 3-Min-Bisection to within a constant factor approximation ratio entails the existence of a constant factor approximation algorithm for the more general and widely-studied problem Min-Bisection, which is defined as above but where $G$ can be arbitrary and where $A$ and $B$ have sizes differing by at most one. Given an arbitrary instance $(G=(V,E),k)$ of size $n$ of 3-Min-Bisection, we now build an instance $(H,\mu,D,\kappa)$ of 2-switched RRP. Moreover, we may assume that $k\leq\frac{n}{3}+46$ as it was proven in [9] that every 3-regular graph has bisection width at most $\frac{n}{3}+46$.
@@
As stated earlier, our static network $H$ is the network $H_{i}\in\mathcal{H}$ where $|H_{i}|=\bar{n}$. We refer to the node set of $H$ as $V^{\prime}$ also and we refer to the subset of nodes within $V^{\prime}$ corresponding to $V$ as $V$ also. Since we are in the 2-swtiched setting, we have one switch $s$ with $2|V^{\prime}|$ ports so that every node of $H$ is adjacent, via switch links, to exactly two ports of the switch. Hence, our switch set is $S=\{s\}$ and our hybrid network is $H(S)$. It is important to note that for any configuration $N$, any node of $H(N)$ can be adjacent to at most $2$ other nodes via dynamic links (as $\Delta_{S}=2$).
@@
Conversely, suppose that $(H,\mu,D,\kappa)$ is a yes-instance of RRP. Let $N$ be a configuration (consisting of a switch matching of $s$) and let $F$ be a collection of flow-paths that witness that the total workload cost for the workload matrix $D$ is at most $\kappa$. Denote by $N$, also, the sub-network of $H(N)$ consisting solely of dynamic links. W.l.o.g. we may assume that every node of $H(N)$ is incident with exactly two dynamic links of $N$ (by adding additional dynamic links that we do not actually use in any flow-path, if necessary).
Claim 1
@@
Suppose that the paths $p_{A}$ and $p_{B}$ have different lengths. From above, the total workload cost of the flow-paths serving the $\beta$-demands is at least $3n^{4}+n(2(\frac{n}{2}+1)^{2}-2n(\frac{n}{2}+1)+n^{2}+n)=3n^{4}+\frac{n^{3}}{2}+n^{2}+2n=\kappa_{\beta}+2n$. Consequently, we must have that $2n\leq\kappa_{1}=\frac{k}{2}+\frac{1}{8}-\frac{1}{4n}+\frac{k}{3n^{2}}$. As $k\leq\frac{n}{3}+46$ (from [9]), we have that $2n\leq\frac{n}{6}+23+\frac{1}{8}-\frac{1}{4n}+\frac{1}{9n}+\frac{46}{3n^{2}}$ which yields a contradiction. Hence, the paths $p_{A}$ and $p_{B}$ each have length $\frac{n}{2}$, with the total workload cost of the flow-paths corresponding to the unit-edges being at most $\kappa_{1}$.
@@
Let $A$ (resp. $B$) be the nodes of $V$ that appear on $p_{A}$ (resp. $p_{B}$). So, $(A,B)$ is a bisection of $G$. Suppose that this bisection has width $k+\epsilon$, for some $\epsilon\geq 1$. So there are $k+\epsilon$ unit-edges whose corresponding work-flows contribute collectively at least $(k+\epsilon)\mu(L+2)\geq(k+1)(\frac{1}{2}+\frac{1}{3n^{2}})=\frac{k}{2}+\frac{1}{2}+\frac{1}{3n^{2}}+\frac{k}{3n^{2}}>\frac{k}{2}+\frac{1}{8}-\frac{1}{4n}+\frac{k}{3n^{2}}=\kappa_{1}$ to the total workload cost, which yields a contradiction. Consequently, $G$ has bisection width at most $k$ and we have that if $(H,\mu,D,\kappa)$ is a yes-instance of RRP then $(G,k)$ is a yes-instance of 3-Min-Bisection. Our result follows as $(H,\mu,D,\kappa)$ can be constructed from $(G,k)$ in time polynomial in $n$.
This result significantly strengthens Theorems 4.1 and 4.2 from [13]: there, RRP($\mathrm{\Delta}_{R}\geq 2,\sigma=0$) is shown to be NP-complete when the static network is an independent set, and the proof does not enable us to restrict the workload matrix $D$ meaningfully. The main weakness of Theorem 3.1 is its reliance on $\mu$ being a polynomial factor smaller than any static link weight. This is actually related to the fact that a connected 2-regular network, as is $G(N)$ when $G$ is an independent set and $\mathrm{\Delta}_{S}=2$, has diameter linear in the number of nodes $n$. A network of maximum degree 3, on the other hand, may have diameter logarithmic in $n$ (e.g., a complete binary tree has this property) and we indeed show NP-completeness of RRP($\mathrm{\Delta}_{S}=3$) when $\mu=\mathrm{\Theta}(\frac{1}{\log n})$.
Theorem 3.2
For any polynomial family of networks =$\mathcal{H}=\{H_{i}:i\geq 0\}$, the problem 3-switched RRP restricted to instances $(H,\mu,D,\kappa)$ satisfying:
@@
  • $\mu\in\mathrm{\Theta}(\frac{1}{\log(n)})$
@@
Proof
As in the case of Theorem 3.1, membership of NP is straightforward. The problem Restricted Exact Cover by 3-Sets (RXC3) is defined as follows:
@@
The problem RXC3 is NP-complete as was proven in Theorem A.1 of [15]. Let $(\mathcal{X},\mathcal{C})$ be an instance of RXC3 of size $n$. We will build an instance $(H,\mu,D,\kappa)$ of 3-switched RRP corresponding to $(\mathcal{X},\mathcal{C})$.
First, we define a weighted graph $D^{\prime}$ so as to describe the workload $D$. Having built $D^{\prime}$, we will orient every undirected edge so that: there is a directed edge $(u,v)$ with weight $w$ if, and only if, there is a node-to-node workload of $w$ from $u$ to $v$ (we still refer to the resulting digraph as $D^{\prime}$). Consequently, we will need one flow-path in our resulting hybrid network corresponding to each directed edge of $D^{\prime}=(V^{\prime},E^{\prime})$. Our construction proceeds in stages.
@@
Let $\bar{n}$ be the size of the network $H_{i}$ where $i$ is the smallest integer such that $N\leq|H_{i}|$ and we fix our static network as $H=H_{i}$. We name a subset of nodes of $H$ as $V^{\prime}$ with the remaining nodes named $U$. $\mathrm{\Delta}_{S}=3$, so we have one switch $s$ that has $3n$ switch ports and every node of $H$ has $3$ external ports; so, every node of $H$ has exactly $3$ switch links to $s$, and $S=\{s\}$. Finally, we define $\mu=\frac{1}{2\lceil\log_{2}(n)\rceil}$ and $\kappa=\mu\alpha|E^{\prime\prime}|+3\mu n(\lceil\log_{2}(n)\rceil+3)$ so as to complete the construction of our instance $(H,\mu,D,\kappa)$ of 3-switched RRP.
Suppose that $(\mathcal{X},\mathcal{C})$ is a yes-instance of RXC3. So, let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be an exact cover of $\mathcal{X}$. For every $(u,v)\in E^{\prime\prime}$, we choose $(u,v)$ to be a dynamic link and ensure that every node of $\{v_{j}:1\leq j\leq 3n,c_{j}\in\mathcal{C}^{\prime}\}$ is joined via a dynamic link to a leaf node of $\{l_{i}:1\leq i\leq n\}$. If $v_{j}$ is such that $c_{j}\in\mathcal{C}^{\prime}$ then denote the unique leaf node to which there is a dynamic link from $v_{j}$ by $\bar{l}_{j}$. Denote the resulting switch matching of $s$ as constituting the configuration $N$. Referring back to our comment on node degrees above, the dynamic links so added must form a matching in $H(N)$. For any workload of weight $\alpha$ in $D$, we choose the corresponding flow-path to consist of the dynamic link joining the two nodes in question. Consequently, for all of the workloads corresponding to directed edges of $E^{\prime\prime}$, the total workload cost of the corresponding flow-paths is $\mu\alpha|E^{\prime\prime}|$. All of the remaining workloads involve the root and the nodes of $\{x_{i}:1\leq i\leq n\}$. Any one of these nodes $x_{i}$ is such that the element $x_{i}$ in a unique clause $c_{j}\in\mathcal{C}^{\prime}$. We choose the flow-path consisting entirely of dynamic links that routes from the root down the tree $T$ to the leaf-node $\bar{l}_{j}$ and on to $v_{j}$ and, through either $u_{j}$ or $w_{j}$, on to $x_{i}$. This flow-path has (graph-theoretic) length $d+3$ and consequently the total workload cost due to flow-paths corresponding to edges of $E^{root}$ is exactly $3\mu n(d+3)=3\mu n(\lceil\log_{2}(n)\rceil+3)$. Hence, the total workload cost due to all flow-paths is $\mu\alpha|E^{\prime\prime}|+3\mu n(\lceil\log_{2}(n)\rceil+3)=\kappa$ and $(H,\mu,D,\kappa)$ is a yes-instance of 3-switched RRP.
Conversely, suppose that $(H,\mu,D,\kappa)$ is a yes-instance of 3-switched RRP and that $N$ is a configuration and $F$ a set of flow-paths witnessing that the total workload cost is at most $\kappa$.
Claim 1
@@
Consider the flow-path corresponding to the directed edge $(r,x_{i})\in E^{root}$. Suppose that the dynamic link $(l_{k},v_{j})$ appears on this flow-path (exactly one such dynamic link does). Note that there are at most $n$ such dynamic links used in the flow-paths of $F$ corresponding to the directed edges of $E^{root}$. Build the set of clauses $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ by including every such $c_{j}$ in $\mathcal{C}^{\prime}$. Consequently, we have a subset $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ of at most $n$ clauses. Moreover, as we can reach any $x_{i}$ from $r$ by a flow-path of $d+3$ dynamic links, every $x_{i}\in\mathcal{X}$ appears in a clause of $\mathcal{C}^{\prime}$; that is, $|\mathcal{C}^{\prime}|=n$, $\mathcal{C}^{\prime}$ is an exact cover for $\mathcal{X}$ and $(\mathcal{X},\mathcal{C})$ is a yes-instance of RXC3. As our instance $(H,\mu,D,\kappa)$ can be constructed from $(\mathcal{X},\mathcal{C})$ in time polynomial in $n$, our result follows.
These results led us to consider the problem 1-switched RRP($\sigma=k$) for $k\geq 0$. By extending Lemma 1 due to [13], we show that this restriction leads to a tractable problem when either $\sigma=0$ or the static network is a complete graph, in contrast with our NP-completeness results. That result makes use of the dynamic link limit $\delta$ recall that when $\delta=1$ every flow-path must contain at most one dynamic link. In combination with the constraint $\sigma=0$, this entails that every flow path consists of either a single dynamic link (and no static links), or of one or more static links (and no dynamic links). The result below is proven via a maximum matching argument in [13].
Lemma 1 ([13, Theorem 3.1])
@@
Theorem 3.3
$1$-switched RRP($\sigma=0$) is in P.
Proof
Note that this proof holds even if we allow variable weights for static and optic links. Because each node $v$ is connected to the switch exactly once, it is not possible that any vertex is incident to two optic links, and hence impossible for there to be any path consisting of two or more optic links. Consequently setting $\delta=1$ introduces no new constraints; in this scenario a flow is permissible with $\sigma=0$ if and only if it permissible with $\sigma=0$ and $\delta=1$. Hence the input instance of 1-switched RRP($\sigma=0,\delta=1$) is a yes-instance if and only the corresponding instance of RRP($\sigma=0,\delta=1$) is a yes-instance, and tractability follows from Lemma 1.
Corollary 1
$1$-switched RRP($\sigma=k$) restricted to instances where the static network $G$ is a complete graph is in P, for any $k\in\mathbb{N}\cup\{\infty\}$.
Proof
@@
  • a single static link from $u$ to $v$ with unit weight.
  • a single dynamic link from $u$ to $v$ with weight $\mu$.
It follows that setting $\sigma=0$ introduces no new constraints, and then by Theorem 3.3 we have tractability.
Corollary 1 rules out the possibility that 1-switched RRP might be NP-complete for any polynomial graph family $\mathcal{H}$ (since such a claim would extend to the family of complete graphs) unless P equals NP. This leaves open the practically relevant case where $\mathrm{\Delta}_{S}=1$ and $\sigma>0$ for specific topologies. We consequently consider the scenario where the static network is a hypercube and the segregation parameter $\sigma=3$.
Theorem 3.4
For any fixed $\mu\in(0,1)$, the problem 1-switched RRP($\sigma=3$) restricted to instances $(H,\mu,D,\kappa)$ satisfying:
  • $H\in\mathcal{Q}$, where $\mathcal{Q}:=\{Q_{d}|d\in\mathbb{N}\}$ is the family of hypercubes